# RAMer: Reconstruction-based Adversarial Model for Multi-party Multi-modal Multi-label Emotion Recognition

### v1

@@ Abstract
Conventional Multi-modal multi-label emotion recognition (MMER) from videos typically assumes full availability of visual, textual, and acoustic modalities. However, real-world multi-party settings often violate this assumption, as non-speakers frequently lack acoustic and textual inputs, leading to a significant degradation in model performance. Existing approaches also tend to unify heterogeneous modalities into a single representation, overlooking each modality’s unique characteristics. To address these challenges, we propose RAMer (Reconstruction-based Adversarial Model for Emotion Recognition), which leverages adversarial learning to refine multi-modal representations by exploring both modality commonality and specificity through reconstructed features enhanced by contrastive learning. RAMer also introduces a personality auxiliary task to complement missing modalities using modality-level attention, improving emotion reasoning. To further strengthen the model’s ability to capture label and modality interdependency, we propose a stack shuffle strategy to enrich correlations between labels and modality-specific features. Experiments on three benchmarks, i.e., MEmoR, CMU-MOSEI, and $M^{3}$, demonstrate that RAMer achieves state-of-the-art performance in dyadic and multi-party MMER scenarios. The code will be publicly available at https://github.com/Sootung/RAMer
@@ 1 Introduction
Emotion recognition from videos is crucial for advancing human-computer interaction and social intelligence. Multi-modal, multi-label emotion recognition (MMER) leverages visual, textual, and acoustic signals to identify multiple emotions (e.g., happy, sad) simultaneously Zhang et al. (2020, 2021a). Traditional MMER methods, as shown in Figure 1(a), typically focus on monologue or dyadic settings, assuming all modalities are fully available. However, real-world conversations often involve multiple participants (i.e., multi-party scenarios) with incomplete modality data for non-speakers who always lack acoustic and textual signals.
Multi-party MMER, a more complex and practical setting, introduces three key challenges. Firstly, handling incomplete modalities is a significant challenge. Most existing approaches Zhang et al. (2021a, 2022); Ge et al. (2023) assume full modality availability and independently encode each modality, ignoring missing data. While some methods Ghosal et al. (2019); Hu et al. (2021) model conversational context by capturing speaker dependencies, they struggle in multi-party scenarios where non-speakers often lack key modalities, leading to poor emotion recognition performance. Secondly, representing diverse modalities effectively remains challenging. Current fusion strategies, such as aggregation-based methods (e.g., concatenation, averaging)Shen et al. (2020) and hybrid approaches Manzoor et al. (2023), project modalities into a shared subspace, often neglecting their unique characteristics and reducing discriminative ability. Recent methods Zhang et al. (2022) attempt to separate modality-specific and shared features but often suffer from information loss due to inadequate handling of inter-modal correlations. Similarly, methods preserving modality-specific information Peng et al. (2023) may overlook cross-modal commonalities, limiting their ability to fully capture inter-modal relationships. Finally, multi-label learning presents challenges in modeling robust label correlations and capturing complex interdependency between modalities and labels. Existing approaches Cisse et al. (2013); Ma et al. (2021) often fail to fully exploit collaborative label relationships. Moreover, emotions vary across modalities, and different emotions rely on distinct modality features, further complicating the task.
@@
As illustrated in Figure 1(b), RAMer addresses the challenge of multi-party MMER by following techniques. To address the challenge of incomplete modalities, we propose an auxiliary task that incorporates external knowledge, such as personality traits, to complement the existing modalities. Leveraging this, we employ modality-level attention mechanisms to capture both inter- and intra-personal features. A reconstruction-based network is utilized to recover the features of any modality by leveraging information from the other modalities.
@@
  • A Novel Model for the Multi-party MMER Problem. We present RAMer, a new model designed to address the Multi-party Multi-modal Multi-label Emotion Recognition problem. RAMer uses adversarial learning to capture both commonality and specificity across multiple modalities, improving emotion recognition even with incomplete modality data.
@@
  • Extensive Experiments. We conduct comprehensive experiments on three benchmarks, i.e., MEmoR, CMU-MOSEI, and $M^{3}$, across various conversation scenarios. Results show that RAMer surpasses existing approaches and achieves state-of-the-art performance in both dyadic and multi-party MMER problems.
@@ 2 Related Work
Emotion recognition has progressed from uni-modal approaches Huang et al. (2021); Saha et al. (2020), which rely on a single modality, to multi-modal methods Mittal et al. (2020); Lv et al. (2021); Zhang et al. (2022) that exploit complementary features across modalities. While uni-modal approaches often face recognition biases Huang et al. (2021), multi-modal learning has gained significant attention, with a key challenge being the effective integration of heterogeneous modalities. Early fusion methods, such as concatenation Ngiam et al. (2011a), tensor fusion Liu et al. (2018a), and averaging Hazirbas et al. (2017), struggle with modality gaps that hinder effective feature alignment. To address this, attention-based methods Ge et al. (2023); Tsai et al. (2019) leverage cross-attention mechanisms to dynamically align features in the latent space, while contrastive learning Chen et al. (2020); Peng et al. (2023) further improves robustness. However, most attention-based methods consolidate modalities into a joint embedding, often overlooking the unique characteristics of each modality.
Figure 2: The framework of RAMer. Given incomplete multi-modal inputs, RAMer first encodes each individual modality through an auxiliary task, then feeds the features into a reconstruction-based adversarial network to extract specificity and commonality. Finally, a stacked shuffle layer is employed to learn enhanced representations.
@@
Multi-label emotion recognition in videos involves assigning multiple emotion labels to a target individual in a video sequence. Early methods treated multi-label classification as independent binary tasks Boutell et al. (2004), but recent advancements explore label correlations using techniques like Adjacency-based Similarity Graph Embedding(ASGE) You et al. (2020), Graph Convolutional Network (GCN) Chen et al. (2019), and multi-task pattern Tsai and Lee (2020) to explore label correlations. Some noteworthy strategies Zhang et al. (2021b, 2022) focus on modeling label-feature correlations through label-specific representations enabled by visual attention Chen et al. (2019) and transformers Zhang et al. (2022). Beyond monologue settings, MMER in conversations has gained interest Poria et al. (2018), using GCN Ghosal et al. (2019) and memory networks Hazarika et al. (2018) to model dynamic speaker interactions. However, multi-party MMER presents significant challenges as it extends beyond recognizing emotions for individual speakers to handling multiple characters with incomplete modalities. Additionally, the influence of individual personalities on label-feature correlations remains underexplored.
@@ 3 Problem Formulation
Each sample is characterized by multiple modalities, including visual ($v$), acoustic ($a$), textual ($t$), and personality traits ($p$).
For each modality $m\in\left\{v,a,t,p\right\}$, the corresponding features are represented as $\left(\bm{\mathcal{X}}^{1},\bm{\mathcal{X}}^{2},\cdots,\bm{\mathcal{X}}^{m}\right)$, where $\bm{\mathcal{X}}^{k}\in\mathbb{R}^{l_{k}\times d_{k}}$ represents the feature space of the $k$-th modality. Here: $l_{k}$ denotes the sequence length, and $d_{k}$ denotes the dimension of the modality.
Let $\mathcal{Y}=\left\{y_{1},y_{2},\cdots,y_{\zeta}\right\}$ represent a label space with $\zeta$ possible emotion labels.
@@ 4.1 Auxiliary Uni-modal Embedding
To extract contextual information from each modality, we employ four independent transformer encoders Vaswani et al. (2017), each dedicated to a specific modality $m$.Each encoder consists of $n_{m}$ identical layers to ensure consistent and deep representation. For real-world multi-party conversation videos with $T$ participants and incomplete modalities, we introduce an optional auxiliary task leveraging personality to complement missing modalities. Specifically, we concatenate personality embedding $\bm{\mathcal{X}}^{p}$ with each modality $\bm{\mathcal{X}}^{m}\in\left\{v,t,a\right\}$ to enrich the feature space. We then apply the scaled dot-product attention to compute inter-person attention across the person dimension within each segment, and intra-person attention along the segment dimension for each individual Shen et al. (2020). This modality-level attention mechanism is designed to enhance the model’s emotion reasoning ability by effectively capturing both interpersonal dynamics and temporal patterns within the data. In this way, we obtain personality enhanced representation $\bm{\mathcal{X}}_{\alpha}^{m}\in\mathbb{R}^{l\times d}$.
@@ 4.2 Reconstruction-based Adversarial Learning
The second component focuses on leveraging multiple modalities by capturing inter-modal commonalities while enhancing the unique characteristics of each modality. To address the limitations of adversarial networks Goodfellow et al. (2014); Zhang et al. (2022), which can result in information loss and difficulty in learning modality-label dependencies in incomplete modality scenarios, we introduce a reconstruction-based approach. This method uses contrastive learning to create modality-independent but label-relevant representations. By reconstructing missing modalities during training, the model ensures more robust performance.
@@
Considering the significance of both specificity and commonality, we employ adversarial training to extract discriminative features. The uni-modal embeddings $\bm{\mathcal{X}}_{\alpha}^{m}$ are fed to three fully connected networks $f_{m}$ to extract specificity $\bm{S}^{m},m\in\left\{v,a,t\right\}$. In parallel, $\bm{\mathcal{X}}_{\alpha}^{m}$ are also passed through a reconstruction network, which is coupled with a contrastive learning network, followed by a generator $G\left(\cdot;\theta_{G}\right)$ to derive the commonality $\bm{C}^{m}$. Both specificity and commonality are then passed through linear layers with softmax activation in the discriminator $D\left(\cdot;\theta_{D}\right)$ that is designed to distinguish which modality the inputs come from. The generator captures commonality representation ${C}^{m}$ by projecting different reconstructed embedding $\bm{\mathcal{X}}_{\gamma}^{m}$ into a shared latent subspace, ensuring distributional alignment across modalities. Consequently, this architecture encourages the generator $G(\cdot;\theta_{G})$ to produce outputs that challenge the discriminator $D\left(\cdot;\theta_{D}\right)$ by obscuring the source modality of the representations $\bm{C}^{m}$. That means the generator and discriminator are trained simultaneously in a game-theoretic scenario, which intends to enhance the robustness of the generated features against modality-specific biases. Both the commonality adversarial loss $\mathcal{L}_{C}$ and the specificity adversarial loss $\mathcal{L}_{S}$ are calculated by cross-entropy loss.
@@
$\vspace{-.1cm}\small\mathcal{L}_{cml}=-\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\sum_{\upsilon=1}^{\zeta}y_{\tau}^{\upsilon}\log\hat{y}_{\tau}^{\upsilon,m}+(1-y_{\tau}^{\upsilon})\log(1-\hat{y}_{\tau}^{\upsilon,m}),$ (1)
where $\hat{y}_{\tau}^{\upsilon,m}$ is predicted with $\bm{C}_{m}$ and $y_{\tau}^{\upsilon}$ is the ground-truth label. In an effort to encode diverse aspects of multi-modal data, an orthogonal loss $\mathcal{L}_{orth}$ is induced to ensure that the encoded subspaces for commonality $\bm{C}^{m}$ and specificity $\bm{S}^{m}$ representations maintain distinctiveness by minimizing overlap.
$\vspace{-.1cm}\small\mathcal{L}_{orth}=-\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\left\|(\bm{C}_{\tau}^{m})^{T}\bm{S}_{\tau}^{m}\right\|_{F}^{2},$ (2)
@@
To reconstruct the features of any modality by leveraging information from the other modalities. We employ a reconstruction network that is composed of modality-specific encoders $\varepsilon^{m}$, decoders $d^{m}$, and a two-level reconstruction process utilizing multi-layer linear networks $g(\cdot)$. Given input $\bm{\mathcal{X}}_{\alpha}^{m}$ from different modality, three encoders $\varepsilon^{m}$ that consist of MLPs are utilized to project $\bm{\bm{\mathcal{X}}}_{\alpha}^{m}$ into latent embedding $\bm{Z}_{\alpha}^{m}$ within the latent space $\mathcal{S}^{z}$. Subsequently, three corresponding decoders $d^{m}$ transform these latent vectors into the decoded vectors $\bm{\mathcal{\widetilde{X}}}_{\alpha}^{m}$. At the first level of reconstruction network, the intrinsic vector $\bm{\widetilde{D}}^{m}$ that derived from contrastive learning network and semantic features $\mathcal{\widetilde{X}}_{\alpha}^{\left\{v,t,a\right\}\backslash m}$ are concatenated to form the input, which is processed to produce $\bm{\mathcal{X}}_{\beta}^{m}$ used for the second-level reconstruction network. Hereby, the reconstruction network can be formulated as,
@@
where $\bm{P}(i)$ is the positive set, $\eta\in\mathbb{R}^{+}$ is a temperature parameter, and $\bm{A}(i)=\bm{\mathcal{Z}}\ \backslash\ {\left\{i\right\}}$.
@@ 4.3 Stack Shuffle for Feature Augmentation
To construct more robust correlations among labels and model the complex interconnections between modalities and labels, we propose a multi-modal feature augmentation strategy that incorporates a stack shuffle mechanism. As shown by Figure LABEL:step3 in Appendix, after obtaining the commonality and specificity representations, we perform sample-wise and modality-wise shuffling processes sequentially on a batch of samples. To strengthen the correlations between labels, we first apply a sample-wise shuffle. The features derived from $\bm{C}$ and $\bm{S}^{m}$ are split into $k$ stacks along the sample dimension, with the top elements of each stack cyclically popped and appended to form new vectors. Next, a modality-wise shuffle is introduced to help the model capture and integrate information across different modalities. For each sample, features are divided into stacks along the modality dimension, and iterative pop-and-append operations are applied. Finally, the shuffled samples $\mathbf{V}$ are used to fine-tune the classifier $c_{\zeta}$ with the binary cross-entropy (BCE) loss.

### v2

@@ Abstract
Conventional Multi-modal multi-label emotion recognition (MMER) assumes complete access to visual, textual, and acoustic modalities. However, real-world multi-party settings often violate this assumption, as non-speakers frequently lack acoustic and textual inputs, leading to a significant degradation in model performance. Existing approaches also tend to unify heterogeneous modalities into a single representation, overlooking each modality’s unique characteristics. To address these challenges, we propose RAMer (Reconstruction-based Adversarial Model for Emotion Recognition), which refines multi-modal representations by not only exploring modality commonality and specificity but crucially by leveraging reconstructed features, enhanced by contrastive learning, to overcome data incompleteness and enrich feature quality. RAMer also introduces a personality auxiliary task to complement missing modalities using modality-level attention, improving emotion reasoning. To further strengthen the model’s ability to capture label and modality interdependency, we propose a stack shuffle strategy to enrich correlations between labels and modality-specific features. Experiments on three benchmarks, i.e., MEmoR, CMU-MOSEI, and $M^{3}$ED, demonstrate that RAMer achieves state-of-the-art performance in dyadic and multi-party MMER scenarios.
@@ 1 Introduction
Emotion recognition from videos is crucial for advancing human-computer interaction and social intelligence. Multi-modal multi-label emotion recognition (MMER) leverages visual, textual, and acoustic signals to identify multiple emotions (e.g., happy, sad) simultaneously Zhang et al. (2020, 2021a). Conventional MMER methods, as shown in Figure 1(a), typically focus on monologue or dyadic settings, assuming all modalities are fully available. However, real-world conversations often involve multiple participants (i.e., multi-party) with incomplete modality data for non-speakers who always lack acoustic and textual signals.
Multi-party MMER, a more complex and practical setting, introduces three key challenges. Firstly, handling incomplete modalities is a significant challenge, which requires robust methods to reconstruct or infer missing information. Most existing works Zhang et al. (2021a, 2022); Ge et al. (2023) assume complete modality access and encode each modality independently, overlooking missing data. While some methods  Ghosal et al. (2019); Hu et al. (2021) leverage speaker-aware context modeling, their performance degrades in multi-party settings where non-speakers often lack critical modalities. Secondly, representing diverse modalities effectively remains challenging, often requiring techniques that can not only integrate disparate information but also reconstruct rich, complete representations from potential modalities. Current fusion strategies, such as aggregation-based methods (e.g., concatenation, averaging)Shen et al. (2020) and hybrid approaches Manzoor et al. (2023), project modalities into a shared subspace, often neglecting their unique characteristics and reducing discriminative ability. Recent methods Zhang et al. (2022) attempt to separate modality-specific and shared features but often suffer from information loss due to inadequate handling of inter-modal correlations. Similarly, methods preserving modality-specific information Peng et al. (2023) may overlook cross-modal commonalities, limiting their ability to fully capture inter-modal relationships. Finally, multi-label learning presents challenges in modeling robust label correlations and capturing complex interdependency between modalities and labels. Existing approaches Cisse et al. (2013); Ma et al. (2021) often fail to fully exploit collaborative label relationships. Moreover, emotions vary across modalities, and different emotions rely on distinct modality features, further complicating the task.
@@
As illustrated in Figure 1(b), RAMer addresses the challenge of multi-party MMER by following techniques. To address the challenge of incomplete modalities, we propose an auxiliary task that incorporates external knowledge, such as personality traits, to complement the existing modalities. Leveraging this, we employ modality-level attention mechanisms to capture both inter- and intra-personal features. A reconstruction-based network is utilized to recover and enrich the features of any modality by leveraging information from the other modalities.
@@
  • A Novel Model for the Multi-party MMER Problem. We present RAMer, a new framework that centrally integrates feature reconstruction within an adversarial learning paradigm. RAMer adeptly captures both commonality and specificity across modalities, crucially utilizing robustly reconstructed features to significantly improve emotion recognition, especially even with incomplete modality data.
@@
  • Extensive Experiments. We conduct comprehensive experiments on three benchmarks, i.e., MEmoR, CMU-MOSEI, and $M^{3}$ED, across various conversation scenarios. Results show that RAMer surpasses existing approaches and achieves state-of-the-art performance in both dyadic and multi-party MMER problems.
@@ 2 Related Work
Emotion recognition has progressed from uni-modal approaches Huang et al. (2021); Saha et al. (2020) to multi-modal methods Lv et al. (2021); Zhang et al. (2022) that exploit complementary features across modalities. While uni-modal approaches often face recognition biases Huang et al. (2021), multi-modal learning has gained significant attention, with a key challenge being the effective integration of heterogeneous modalities. Early fusion methods, such as concatenation Ngiam et al. (2011a), tensor fusion Liu et al. (2018a), and averaging Hazirbas et al. (2017), struggle with modality gaps that hinder effective feature alignment. To address this, attention-based methods Ge et al. (2023); Tsai et al. (2019) leverage cross-attention mechanisms to dynamically align features in the latent space, while contrastive learning Chen et al. (2020); Peng et al. (2023) further improves robustness. However, most attention-based methods merge modalities into a joint embedding, often overlooking modality-specific features.
Figure 2: The framework of RAMer. Given incomplete multi-modal inputs, RAMer first encodes each individual modality through an auxiliary task, then feeds the features into a reconstruction-based adversarial network to extract specificity and commonality. Finally, a stack shuffle strategy is employed to learn enhanced representations.
@@
MMER involves assigning multiple emotion labels to a target person in the video. Early methods treated multi-label classification as independent binary tasks Boutell et al. (2004), but advancements explore label correlations using techniques like Adjacency-based Similarity Graph Embedding(ASGE) You et al. (2020), Graph Convolutional Network (GCN) Chen et al. (2019), and multi-task pattern Tsai and Lee (2020) to explore label correlations. Some noteworthy strategies Zhang et al. (2021b, 2022) focus on modeling label-feature correlations through label-specific representations enabled by visual attention Chen et al. (2019) and transformers Zhang et al. (2022). Beyond monologue settings, MMER in conversations has gained interest, using GCN Ghosal et al. (2019) and memory networks Hazarika et al. (2018) to model dynamic speaker interactions. However, multi-party MMER is especially challenging, as it requires recognizing emotions for multiple speakers with incomplete modalities. Additionally, the influence of individual personalities on label-feature correlations remains underexplored.
@@ 3 Problem Formulation
Each sample involves modalities such as visual ($v$), acoustic ($a$), textual ($t$), and personality traits ($p$).
For each modality $m\in\left\{v,a,t,p\right\}$, the corresponding features are represented as $\left(\bm{\mathcal{X}}^{1},\bm{\mathcal{X}}^{2},\cdots,\bm{\mathcal{X}}^{m}\right)$, where $\bm{\mathcal{X}}^{k}\in\mathbb{R}^{l_{k}\times d_{k}}$ represents the feature space of the $k$-th modality. Here: $l_{k}$ denotes the sequence length, and $d_{k}$ denotes the dimension of the modality. Let $\mathcal{Y}=\left\{y_{1},y_{2},\cdots,y_{\zeta}\right\}$ represent a label space with $\zeta$ possible emotion labels.
@@ 4.1 Auxiliary Uni-modal Embedding
To extract contextual information from each modality, we employ four independent transformer encoders Vaswani et al. (2017), each dedicated to a specific modality $m$.Each encoder consists of $n_{m}$ identical layers to ensure consistent and deep representation. For multi-party conversation videos with $T$ participants with incomplete modalities, we introduce an optional auxiliary task leveraging personality to complement missing modalities. Specifically, we concatenate personality embedding $\bm{\mathcal{X}}^{p}$ with each modality $\bm{\mathcal{X}}^{m}\in\left\{v,t,a\right\}$ to enrich the feature space. We then apply the scaled dot-product attention to compute inter-person attention across the person dimension within each segment, and intra-person attention along the segment dimension for each individual Shen et al. (2020). This modality-level attention mechanism is designed to enhance the model’s emotion reasoning ability by effectively capturing both interpersonal dynamics and temporal patterns within the data. In this way, we obtain personality enhanced representation $\bm{\mathcal{X}}_{\alpha}^{m}\in\mathbb{R}^{l\times d}$.
@@ 4.2 Reconstruction-based Adversarial Learning
The second component leverages multiple modalities by capturing inter-modal commonalities while preserving modality-specific features. To address the limitations of adversarial networks Goodfellow et al. (2014); Zhang et al. (2022), which can result in information loss and difficulty in learning modality-label dependencies, we introduce a reconstruction-based approach. It employs contrastive learning to learn modality-independent but label-relevant representations while reconstructing missing modalities during training to enhance robustness.
@@
To balance modality specificity and commonality, we adopt adversarial training to extract discriminative features. The uni-modal embeddings $\bm{\mathcal{X}}_{\alpha}^{m}$ are fed to three fully connected networks $f_{m}$ to extract specificity $\bm{S}^{m},m\in\left\{v,a,t\right\}$. In parallel, $\bm{\mathcal{X}}_{\alpha}^{m}$ are also passed through a reconstruction network, which is coupled with a contrastive learning network, followed by a generator $G\left(\cdot;\theta_{G}\right)$ to derive the commonality $\bm{C}^{m}$. Both specificity and commonality are then passed through linear layers with softmax activation in the discriminator $D\left(\cdot;\theta_{D}\right)$ that is designed to distinguish which modality the inputs come from. The generator captures commonality ${C}^{m}$ by projecting different reconstructed embedding $\bm{\mathcal{X}}_{\gamma}^{m}$ into a shared latent subspace, ensuring distributional alignment across modalities. Consequently, this architecture encourages the generator $G(\cdot;\theta_{G})$ to produce outputs that challenge the discriminator $D\left(\cdot;\theta_{D}\right)$ by obscuring the source modality of $\bm{C}^{m}$. The generator and discriminator are jointly trained in a game-theoretic setup to enhance feature robustness against modality-specific biases. Both the commonality adversarial loss $\mathcal{L}_{C}$ and the specificity adversarial loss $\mathcal{L}_{S}$ are calculated by cross-entropy loss as,
$\small\mathcal{L}_{C}=-\frac{1}{N}\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\Big{(}U^{m}\log\big{(}D(C_{\tau}^{m};\theta_{D})\big{)}\Big{)},$ (1)
$\small\mathcal{L}_{S}=-\frac{1}{N}\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\Big{(}U^{m}\log\big{(}D(S_{\tau}^{m};\theta_{D})\big{)}\Big{)},$ (2)
where $U\in\left\{U^{v},U^{t},U^{a}\right\}$ represents the ground truth label corresponding to the discriminator’s input.
@@
$\small\mathcal{L}_{cml}=-\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\sum_{\upsilon=1}^{\zeta}y_{\tau}^{\upsilon}\log\hat{y}_{\tau}^{\upsilon,m}+(1-y_{\tau}^{\upsilon})\log(1-\hat{y}_{\tau}^{\upsilon,m}),$ (3)
where $\hat{y}_{\tau}^{\upsilon,m}$ is predicted with $\bm{C}_{m}$ and $y_{\tau}^{\upsilon}$ is the ground-truth label. To encode diverse aspects of multi-modal data, we introduce an orthogonal loss $\mathcal{L}_{orth}$ that encourages the commonality $\bm{C}^{m}$ and specificity $\bm{S}^{m}$ subspaces to remain distinct by minimizing their overlap.
$\small\mathcal{L}_{orth}=-\sum_{m\in\{v,t,a\}}\sum_{\tau=1}^{N}\left\|(\bm{C}_{\tau}^{m})^{T}\bm{S}_{\tau}^{m}\right\|_{F}^{2},$ (4)
@@
[…]
@@
[…]
@@ 4.3 Stack Shuffle for Feature Augmentation
To construct more robust correlations among labels and model the complex interconnections between modalities and labels, we propose a multi-modal feature augmentation strategy that incorporates a stack shuffle mechanism. After obtaining the commonality and specificity representations, we perform sample-wise and modality-wise shuffling processes sequentially on a batch of samples. To strengthen the correlations between labels, we first apply a sample-wise shuffle. The features derived from $\bm{C}^{m}$ and $\bm{S}^{m}$ are split into $k$ stacks along the sample dimension, with the top elements of each stack cyclically popped and appended to form new vectors. Next, a modality-wise shuffle is introduced to help the model capture and integrate information across different modalities. For each sample, features are divided into stacks along the modality dimension, and iterative pop-and-append operations are applied. Finally, the shuffled samples $\mathbf{V}$ are used to fine-tune the classifier $c_{\zeta}$ with the binary cross-entropy (BCE) loss.